OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie?

OpenStudy (anonymous):

Write an expression: Let n represent the number of nights: 2.00 + 0.50n = 5.00 Now, isolate the variable by subtracting 2.00 from the other side 0.50n = 5.00 - 2.00 0.50n = 3.00 Divide both sides by 0.50 to isolate "n" 3.00 / 0.50 = 6 Since n represents the number of nights, Laura rented the movie for 6 nights. Hope this helped!

OpenStudy (anonymous):

Thank you so much :)